### v4

# Towards the Causal Complete Cause of Multi-Modal Representation Learning

Abstract
Multi-Modal Learning (MML) aims to learn effective representations across modalities for accurate predictions. Existing methods typically focus on modality consistency and specificity to learn effective representations. However, from a causal perspective, they may lead to representations that contain insufficient and unnecessary information. To address this, we propose that effective MML representations should be causally sufficient and necessary. Considering practical issues like spurious correlations and modality conflicts, we relax the exogeneity and monotonicity assumptions prevalent in prior works and explore the concepts specific to MML, i.e., Causal Complete Cause ($C^{3}$). We begin by defining $C^{3}$, which quantifies the probability of representations being causally sufficient and necessary. We then discuss the identifiability of $C^{3}$ and introduce an instrumental variable to support identifying $C^{3}$ with non-exogeneity and non-monotonicity. Building on this, we conduct the $C^{3}$ measurement, i.e., $C^{3}$ risk. We propose a twin network to estimate it through (i) the real-world branch: utilizing the instrumental variable for sufficiency, and (ii) the hypothetical-world branch: applying gradient-based counterfactual modeling for necessity. Theoretical analyses confirm its reliability. Based on these results, we propose $C^{3}$ Regularization, a plug-and-play method that enforces the causal completeness of the learned representations by minimizing $C^{3}$ risk. Extensive experiments demonstrate its effectiveness.
Machine Learning, ICML
1 Introduction
The initial inspiration for artificial intelligence is to imitate human perceptions which are based on different modalities (Jordan & Mitchell, 2015; Mahesh, 2020), e.g., sight, sound, and touch. Each modality provides unique information with distinct properties (Baltrušaitis et al., 2018; Wang et al., 2023c). Multi-modal learning (MML) (Huang et al., 2021a; Ge et al., 2023) has emerged as a promising approach to emulate human perception. It aims to learn good representations from multiple modalities for accurate decisions.
In this context, a key question arises: what defines a “good” MML representation? Existing methods typically define good MML representations from two perspectives: modality consistency and modality specificity (Zhang et al., 2023a; Ge et al., 2023; Radford et al., 2021; Fan et al., 2024; Dong et al., 2024). The former emphasizes extracting modality-shared semantics related to primary events within the MML task. This type of methods (Liang et al., 2022; Xia et al., 2024; Abdollahzadeh et al., 2021) aims to obtain unified representations by mapping features from different modalities into a common embedding space. In contrast, the second perspective believes modality specificity captures the distinct statistical properties of each modality, reflecting different aspects of the primary events. This type of methods (Dong et al., 2024; Yang et al., 2022; Frost et al., 2015; Zhou et al., 2021) decomposes features within each modality into modality-specific and modality-shared components, learning all shared components while applying distance constraints on modality-specific features to enhance diversity.
Figure 1: Example of causal sufficiency and necessity in “duck” classification task (See Section 2 for more analyses).
However, from a causal perspective (Ahuja et al., 2020; Koyama & Yamaguchi, 2020), these methods may result in the learned representations being insufficient or unnecessary. Specifically, sufficiency indicates that the use of the representations will establish the label, while necessity indicates that the label becomes incorrect when the representations are absent (Pearl, 2009). Figure 1 provides an example. If the MML model only focuses on causal sufficiency, it will lose important modality-specific semantics, affecting generalization; if the model only focuses on causal necessity, the decisions may be made incorrectly based on the background, affecting discriminability. Correspondingly, considering the aforementioned two types of MML methods: (i) those focused on modality consistency may cause the model to overlook important modality-specific semantics; (ii) those focused on modality specificity may incorrectly capture irrelevant information. Thus, existing MML methods (Dong et al., 2024; Lu et al., 2022) without causal constraints may fail to satisfy sufficiency and necessity, affecting model performance. The experiments in Section 6.2 further prove this (Figure 4 and Table 1): (i) the representations learned by existing methods have much lower correlation scores with sufficient and necessary causes than the proposed method with specific constraints; (ii) after constraining the learned representations, the performance of existing MML methods are significantly improved. The analyses in Section 4 also emphasize the importance of causal sufficiency and necessity. Thus, inspired by (Pearl, 2009; Yang et al., 2024), we propose that a good MML representation must with both causal sufficiency and necessity, i.e., being causal complete.
Noticeably, in practical MML applications, constraining the causal sufficiency and necessity of learned representations remains challenging. Specifically, the related discussion in existing works (Yang et al., 2024; Pearl, 2009) are based on the exogeneity and monotonicity assumptions (Definition D.1 and Definition D.2 in Appendix D.4). Exogeneity refers to the scenario where the influence of the external intervention on the conditional distributions is negligible when the causal representation variable is exogenous relative to label variable, while monotonicity illustrates the consistent, unidirectional effect on the label variable of representation variable. However, in practice, non-trivial corner cases often arise: the inseparability of MML semantics leads to spurious correlations (Figure 2), and cross-modal conflicts combined with high-dimensional nonlinear interactions (Huang et al., 2021b; Wang et al., 2023d) undermine monotonicity. These challenges render the conditions of traditional causal sufficiency and necessity discussion inapplicable.
To address this, in this paper, we relax the above assumptions to explore the causal sufficiency and necessity in MML, ensuring the quality of the learned representation. Firstly, we propose the definition of causal sufficiency and necessity, i.e., causal complete causes ($C^{3}$), for MML based on (Pearl, 2009). It reflects the probability that the learned representation is causally complete by estimating the probability of label change after an intervention on the representation given two conditions of observations. One condition is for sufficiency, and another is for necessity. Then, we analyze the identifiability of $C^{3}$, which allows us to quantify $C^{3}$ using the observable data in practice. Unlike previous studies, we propose an instrumental variable to ensure estimating the segmented effect of $C^{3}$ without spurious correlations, thereby relaxing the assumptions of exogeneity and monotonicity. Based on this, we propose the measurement of $C^{3}$ using the twin network, i.e., $C^{3}$ risk, where a low $C^{3}$ risk means that the learned representations are causally complete with high confidence. The challenge of $C^{3}$ measurement lies in eliminating the spurious correlations in the sufficiency evaluation and modeling the counterfactual data required for the necessity evaluation. The proposed twin network addresses this through (i) the real-world branch: using the proposed instrumental variable to eliminate spurious correlations in the representation, and (ii) the hypothetical-world branch: using provable gradient-based adjustments to model the counterfactuals. Through theoretical analyses, we prove the reliability of the twin network and provide the performance guarantee of the $C^{3}$ risk. Based on these theoretical results, we propose Causal Complete Cause Regularization ($C^{3}$), a plug-and-play method to learn causal complete representations by constraining their $C^{3}$ risks.
The main contributions are as follows: (i) We propose the definition, identifiability, and measurement of the causal complete cause ($C^{3}$) concept for MML without the assumptions of exogeneity and monotonicity. These results provide an effective way, i.e., $C^{3}$ risk, to estimate the sufficiency and necessity of the learned representation for robust and accurate MML (Section 3). (ii) We theoretically demonstrate the effectiveness and reliability of the proposed $C^{3}$ risk. Inspired by this, we propose $C^{3}$, which can be applied to any MML model to learn causal complete representations with low $C^{3}$ risk ((Sections 4 and 5). (iii) Finally, we conduct extensive experiments on various datasets that prove the effectiveness and robustness of $C^{3}$ (Section 6).
2 Problem Analysis
Problem Settings
Let $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{Z}$ represent the input space, label space, and latent space, respectively. Consider a MML task involving data $(x,y)\in\mathcal{X}\times\mathcal{Y}$, where $x\in\mathcal{X}$ denotes the sample and $y\in\mathcal{Y}$ denotes the label. All the data are sampled from the same joint distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$. The training dataset is defined as $\mathcal{D}_{tr}=\{(x_{i},y_{i})\}_{i=1}^{N}$, and the testing dataset is defined as $\mathcal{D}_{te}$. Here, each sample $x_{i}=\{x^{(1)}_{i},\dots,x^{(K)}_{i}\}$ contains $K$ modalities. The objective of MML is to obtain a robust model $f_{\theta}=\mathcal{W}\circ g$ that performs well on unseen test dataset $\mathcal{D}_{te}$. Here, the $g:\mathcal{X}\rightarrow\mathcal{Z}$ is the feature extractor and $\mathcal{W}:\mathcal{Z}\rightarrow\mathcal{Y}$ is the classifier. The specific implementation is illustrated in Appendix G, and the table of notations is shown in Table 4.
From the perspective of data generation (Suter et al., 2019; Hu et al., 2022b), each sample is composed of various generating factors, e.g., color, shape, etc. In the context of MML, each sample contains generating factors from all modalities. From a causal perspective (Pearl, 2009), these generating factors can be divided into task-related factors that come from the label, i.e., causal factors $\text{F}_{c}$, and task-independent factors, i.e., non-causal factors $\text{F}_{s}$. The key to accurate prediction lies in extracting the causal generating factors $\text{F}_{c}$ that are related to labels in the latent space (Deshpande et al., 2022). Thus, the goal of $f_{\theta}$ can be re-established as learning representations that contain all causal generating factors $\text{F}_{c}$ from the data without $\text{F}_{s}$ for accurate prediction.
Example of Causal Sufficiency and Necessity
As shown in Figure 1, assume that the training data is with three modalities of image, text, and audio, and all the modalities have the feature “duck paws” when a “duck” exists. The goal of this MML task is to classify the “duck”. Then, the models that learn representations based on consistency will capture the features of “duck paws”, and can establish the label “duck” based on the learned representations. However, the learned model may also make errors in another “duck” scenario where the MML samples do not contain “duck paws”, e.g., “duck swim on the lake” (upper right in Figure 1). This suggests that the learned representation contains sufficient but unnecessary causes since the label “duck” can be predicted using the current representation, but may not work in another scenario. Similarly, there are some necessary but insufficient representations, e.g., “duck” must with “wings”, but “wings” may also correspond to another label, e.g., “bird” (lower of Figure 1). For sufficient and necessary causes, they ensure that the learned representations not only reflect the features target the “duck” label, e.g., “duck paws”, but also the features where the “duck” label must have, e.g., “wings”. Thus, good representations must have both causal sufficiency and necessity, i.e., causal complete causes, which is the goal we explore in this study.
Figure 2: Structural Causal Model (SCM) for MML. Left: causal generating mechanism, Right: the learning process.
Causal Analysis of MML
We conduct a Structural Causal Model (SCM) for MML based on causal generating mechanism  (Suter et al., 2019; Deshpande et al., 2022), as shown in Figure 2 left. In this SCM, Y and X denote the label variable and corresponding generated data variable in the MML task. $\text{F}_{c}$ and $\text{F}_{s}$ represent the distinct sets of generating factors that are causally and non-causally related to Y. Each generating factor corresponds to a semantic of the MML task, e.g., color, shape, background, modality indicator, etc. Since both $\text{F}_{c}$ and $\text{F}_{s}$ represent high-level knowledge of the data but only $\text{F}_{c}$ is causally related to Y, we could naturally define the MML task label variable Y as the cause of the $\text{F}_{c}$, i.e., ${\text{Y}}\rightarrow\text{F}_{c}$. There is no connection between $\text{F}_{s}$ and Y. Following (Hu et al., 2022a; Deshpande et al., 2022), the sample in each modality is generated simultaneously using all the generating factors, including factors caused from the label and unobservable variables, e.g., environmental effects. The unobservable variable results in the generating factors $\text{F}_{s}$, which is non-causally related to Y. Then, we get both $\text{F}_{c}\rightarrow\text{X}$ and $\text{F}_{s}\rightarrow\text{X}$. Thus, we obtain Figure 2 left. Based on this, we further construct a SCM to discuss the learning process of MML, as shown in Figure 2 right. It can be viewed as the inverse process of the causal generating mechanism. Based on Figure 2 left, an ideal MML model should only utilize causal generating factors $\text{F}_{c}$ and be invariant to any intervention on non-causal generating factors $\text{F}_{s}$. However, in practice, $\text{F}_{c}$ and $\text{F}_{s}$ in the learned representation may be coupled (Dong et al., 2024; Li et al., 2023). This leads to the model potentially learning based on non-causal $\text{F}_{s}$. There exist spurious correlations between $\text{F}_{s}$ and Y, i.e., we get the additional $\text{F}_{s}\rightarrow\text{Y}$ in Figure 2 right.
The presence of spurious correlations leads to inaccuracies in MML. Specifically, the learned MML representations may fall into four scenarios: (i) sufficient and necessary: contains all $\text{F}_{c}$; (ii) necessary but insufficient: contains part of $\text{F}_{c}$; (iii) sufficient but unnecessary: contains all $\text{F}_{c}$ with $\text{F}_{s}$; (iv) insufficient and unnecessary: only contains $\text{F}_{s}$. See Appendix D and G for more analyses. To ensure accurate and robust MML, in this paper, we aim to propose a methodology that constrains the model to learn representations only contain causally sufficient and necessary information.
3 Causal Complete Cause
To access the learning of causal sufficient and necessary representations, in this section, we first provide the definition of causal sufficiency and necessity in MML, i.e., Causal Complete Cause ($C^{3}$). Next, we discuss the identifiability of $C^{3}$, which ensures the quantification of $C^{3}$ through observable data in practice, even in corner cases (non-monotonicity and non-exogeneity). Finally, we give the measurement of $C^{3}$, i.e., $C^{3}$ risk to measure the probability of whether the learned MML representation is causally complete.
3.1 Definition of $C^{3}$
To learn representations of causal variable $\text{F}_{c}$ for MML that is with both causal sufficiency and necessity, based on (Pearl, 2009; Yang et al., 2024), we introduce the concept of the probability of Causal Complete Cause ($C^{3}$) as follows:
Definition 3.1 (Probability of Causal Complete Cause ($C^{3}$)).
Denote that the data and corresponding label variables of the given multi-modal data distribution are X and Y, while the variable of the learned MML representation is Z. Let the specific implementations of representation variable Z as $c$ and $\bar{c}$, where $c$ denotes the implementation that results in the accurate label prediction $\text{Y}=y$, and $\bar{c}\not=c$ denotes the implementation resulting in $\text{Y}\not=y$. The probability that Z is the causal complete cause of Y can be defined as:
[…]
where $P(\text{Y}_{do(\text{Z}=c)}=y\mid\text{Z}=\bar{c},\text{Y}\not=y)$ denotes the probability of $\text{Y}=y$ when force Z to be another specific implementation $c$ via do-operator $do(\text{Z}=c)$, when giving observations $\text{Z}=\bar{c}$ and $\text{Y}\not=y$ with probability $P(\text{Z}=\bar{c},\text{Y}\not=y)$. Similarly, the second term corresponds to the case that where the observations are $\text{Z}=c$ and $\text{Y}=y$, the probability of Y becomes incorrect when force $\text{Z}=\bar{c}$.
Definition 3.1 indicates that when Z with a high $C^{3}$ score, it has a high probability of being the causal complete cause of Y. The two terms in Eq.1 correspond to sufficiency and necessity, respectively. Specifically, the sufficiency term $P(\text{Y}_{do(\text{Z}=c)}=y\mid\text{Z}=\bar{c},\text{Y}\not=y)$ means that even when $\text{Z}=\bar{c}$ and $\text{Y}\neq y$, intervening to set $\text{Z}=c$ results in a high probability of $\text{Y}=y$, indicating that $\text{Z}=c$ has a sufficient causal effect on the occurrence of $\text{Y}=y$. Correspondingly, for necessity term $P(\text{Y}_{do(\text{Z}=\bar{c})}\not=y\mid\text{Z}=c,\text{Y}=y)$, even when $\text{Z}=c$ and $\text{Y}=y$, intervening to set $\text{Z}=\bar{c}$ results in $\text{Y}\neq y$, indicating that $\text{Z}=c$ is necessary for the occurrence of $\text{Y}=y$. Under this condition, the learned representation can be divided into sufficient but unnecessary causes, necessary but insufficient causes, sufficient and necessary causes, and insufficient and unnecessary causes (Appendix D for detailed analyses). We aim to constrain the $C^{3}$ score of the learned representation, i.e., extract sufficient and necessary causes, to achieve robust and accurate MML.
3.2 Identifiability of $C^{3}$
Since it is difficult to obtain all samples in the multi-modal data distribution, especially in real systems, e.g., the counterfactual data in the definition of $C^{3}$ is difficult to obtain (Kusner et al., 2017; Morgan & Winship, 2015), calculating the probability of $C^{3}$ is still is a challenging issue. To access the calculation of $C^{3}$ based on observable MML data, we discuss the identifiability of $C^{3}$ in this section.
Identifiability refers to the ability to uniquely infer causal effects from observable data under given assumptions (Pearl, 2009). For MML settings, it means the causal factors within the representation $Z$ can uniquely determine the label $Y$ from the sample $X$. To ensure reliable estimation, we constrain the $C^{3}$ score of the learned representation $Z$ extracted from $X$, ensuring its identifiability. In previous studies (Tian & Pearl, 2002; Yang et al., 2024), identifying causal probabilities typically assumes that statistical data are derived under exogeneity and monotonicity conditions (Definition D.1 and Definition D.2 in Appendix D.4). Exogeneity refers to where the external intervention’s impact on the conditional distribution $P(\text{Y}|\text{Z})$ is negligible when the variable Z is exogenous to Y. This ensures the learned Z to recognize $\text{F}_{c}$ without being affected by $\text{F}_{s}$. Monotonicity, on the other hand, describes the consistent, unidirectional effect of the representation Z on Y. This ensures that the impact of the learned representation Z on the labeling decision will not be reversed as other factors change. However, in practice, non-trivial cases often arise: the inseparability of generating factors in the latent space leads to spurious correlations, i.e., the learned representations contain non-causal $\text{F}_{s}$ (Figure 2); Meanwhile, modality conflicts and the high-dimensional nonlinear interactions (Wang et al., 2023d; Huang et al., 2021b) undermine monotonicity. These challenges render traditional identifiability conditions inapplicable. To address this, we relax the above assumptions and explore the identifiability of $C^{3}$, accessing the calculation of $C^{3}$ score.
Consider the effects of spurious correlations, i.e., $P(\text{Y}_{do(\text{Z}=c)})\neq P(\text{Y}\mid\text{Z}=c)$, we first relax the exogeneity assumption. Based on Definition 3.1, we extend Theorem 9.2.15 in (Pearl, 2009) to MML for $C^{3}$, obtaining:
Theorem 3.2 (Identifiability under Non-Exogeneity).
Given the MML model $f_{\theta}$, where the label variable Y is influenced by the causal factors $\text{F}_{c}$ and non-causal factors $\text{F}_{s}$. Assume $f_{\theta}$ satisfy the Positive Markovian assumption (Tian & Pearl, 2002), the probabilities of $C^{3}$ ($P(\text{Y}_{do(\text{Z}=c)}=y,\text{Y}_{do(\text{Z}=\bar{c})}\neq y)$) is identifiable and can be estimated via:
$C^{3}(\text{Z})=P(\text{Y}_{do(\text{Z}=c)})-P(\text{Y}_{do(\text{Z}=\bar{c})})$ (2)
where $Y$ is monotonic relative to Z with $\text{Y}_{do(\text{Z}=c)}=\bar{y}\wedge\text{Y}_{do(\text{Z}=c)}=y$ is false or $\text{Y}_{do(\text{Z}=\bar{c})}=y\wedge\text{Y}_{do(\text{Z}=c)}=\bar{y}$ is false, and the model $f_{\theta}$ satisfies local invertibility.
Theorem 3.2 establishes that, under the monotonicity condition, $C^{3}$ can be estimated using observable multi-modal data, thereby quantifying $C^{3}$. Specifically, by modeling the effects of spurious correlations ($\text{F}_{s}$ on Y) as non-causal pathways in the SCM, interventions with the $do$-operator, such as $P(\text{Y}_{do(\text{Z}=c)})$ and $P(\text{Y}_{do(\text{Z}=\bar{c})})$, facilitates reliable causal effect estimation. For counterfactual terms ($\text{Z}=\bar{c}$) involved in the necessity probability, the model $f_{\theta}$ imposes locally reversible constraints to ensure identifiability, as discussed in (Galles & Pearl, 1998). Under these conditions, the probability of causal sufficiency $C^{3}_{su}$ and the probability of causal necessity $C^{3}_{ne}$ can be estimated as follows:
$C^{3}_{su}(\text{Z})=\frac{P(\text{Y}_{do(\text{Z}=c)})-P(\text{Y}_{do(\text{Z}=\bar{c})})}{1-P(\text{Z}=c)}$
$C^{3}_{ne}(\text{Z})=\frac{P(\text{Y}_{do(\text{Z}=c)})-P(\text{Y}_{do(\text{Z}=\bar{c})})}{P(\text{Z}=c)}$
Next, according to the second paragraph of Subsection 3.2, considering the cross-modal conflicts with high-dimensional nonlinear interactions in practice, we relax the assumptions of monotonicity. The key idea lies in introducing an instrumental variable $V$ (Caner & Hansen, 2004) to estimate the piecewise effects of conditional distributions within the $C^{3}$ definition. Meanwhile, $V$ satisfies $V\perp\!\!\!\perp\text{F}_{s}$ with a controllable impact on $\text{F}_{c}$, i.e., $V\to\text{F}_{c}\to\text{Y}$. Then, we have:
Theorem 3.3 (Identifiability under Non-Monotonicity and Non-Exogeneity).
[…]
$\begin{split}C^{3}(\text{Z})&=\int_{v}\big{[}P(\text{Y}=y\mid\text{Z}=c,V=v)\\ &-P(\text{Y}=y\mid\text{Z}=\bar{c},V=v)\big{]}P(V=v)\,\text{d}v.\end{split}$ (5)
Theorem 3.3 posits that the estimation of $C^{3}$ is achievable without the dependence on exogeneity and monotonicity, enabling the quantification of $C^{3}$ in the absence of counterfactual data. The detailed discussion is provided in Appendix D.4, and the proofs of the above theorems are provided in Appendix A.1 and A.2, respectively. According to this theorem, the practical estimation of $C^{3}$ hinges on how the instrumental variable $V$ is modeled and how the sufficiency and necessity terms of $C^{3}$ are constrained under $V$, also the main focus in the next subsection, i.e., measurement of $C^{3}$.
3.3 Measurement of $C^{3}$
Based on Definition 3.1 and Theorem 3.3, we provide the measurement of $C^{3}$ in this section, i.e., $C^{3}$ risk, to estimate the $C^{3}$ score of the representation distribution $P(\text{Z}|\text{X}=x)$ inferred from X on the multi-modal data distribution $P_{XY}$. When the learned representation obtains less necessary and sufficient information, the $C^{3}$ risk will be higher. Specifically, we first model the instrumental variable $V$ based on Theorem 3.3. It aims to make the MML model learn representations Z that only contain the causal factors $\text{F}_{c}$ for decision-making while ensuring the identifiability of the $C^{3}$ score assessment (Theorem 3.4). Next, based on Definition 3.1, we provide the $C^{3}$ measurement of the learned Z, i.e., $C^{3}$ risk (Theorem 3.5). We employ a twin network (Figure 5) with a real-world branch (extract causal representation using the proposed instrumental variable) and a hypothetical-world branch (modeling counterfactual data with gradient-based adjustments), thereby modeling the sufficiency and necessity terms. Appendix D.5 establishes the reliability of the proposed twin network.
Firstly, as shown in Section 2, we establish the model $f_{\theta}$ using the feature extractor $g$ and the linear classifier $\mathcal{W}:\mathbb{R}^{d}\rightarrow\mathcal{Y}$ on the overall representation $\text{Z}=g(\text{X})$ ($\text{Z}=\Xi_{c}\text{F}_{c}+\Xi_{s}\text{F}_{s}$) to obtain the label $\text{Y}={\rm sign}(\mathcal{W}^{\top}g(\text{X}))$. To infer $\text{F}_{c}$ from observable data $x\sim\mathcal{X}$, we use the instrumental variable $V$. Here, we focus on how $V$ is modeled to constrain Z only contains $\text{F}_{c}$ for decision-making while satisfying the conditions in Theorem 3.3. We propose:
Theorem 3.4 (Instrumental Variable $V$ in MML).
For multimodal input $X\in\mathbb{R}^{K\times d}$, representing $K$ modalities each with a feature dimension of $d$, instrumental variable $V$ is introduced to extract causal generating factors $\text{F}_{c}$ while reducing the influence of non-causal factors $\text{F}_{s}$. $V$ is generated from $X$ using self-attention mechanism with:
$\begin{split}V=\sum_{j=1}^{K}\frac{\exp(s_{ij})}{\sum_{k=1}^{K}\exp(s_{ik})}X_{j}W_{V},\\ \text{where}\quad s_{ij}=\frac{q_{i}^{\top}k_{j}}{\sqrt{d}}-\alpha\|q_{i}-k_{j}\|^{2}.\end{split}$ (6)
where $q_{i}=X_{i}W_{Q}$ and $k_{j}=X_{j}W_{K}$ are the query and key of $i$-th and $j$-th modalities’ features, $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ are linear projection matrices, mapping query, key and value respectively. Then, $V$ satisfies (i) $P(\text{Z}\mid V)\neq P(\text{Z})$; (ii) $P(\text{F}_{s}\mid V)=P(\text{F}_{s})$; and (iii) $P(\text{Y}\mid\text{Z},V)=P(\text{Y}\mid\text{F}_{c})$.
In Theorem 3.4, we propose using the self-attention mechanism to model the instrumental variable $V$ for $C^{3}$. Specifically, the alignment scores $s_{ij}$ is to capture inter-modal interactions for label. Then, by strengthening the correlation between $V$ and Z, the learned representation Z is guided to approach $\text{F}_{c}$ while weakening the correlation with $\text{F}_{s}$. Furthermore, consider inter-modal misalignment, $\|q_{i}-k_{j}\|^{2}$ with weight $\alpha$ is introduced as a distance-based adjustment to diminish the impact of modalities with large feature distances on attention weights. It satisfies three conditions: (i) Relevance: $V$ manipulate $\text{F}_{c}$; (ii) Independence: $V$ is conditionally independent from $\text{F}_{s}$; and (iii) Exclusion: the influence of $V$ on Y is completely indirectly realized through $\text{F}_{c}$. The detailed proofs are provided in Appendix A.3.
Next, based on these results, we provide the measurement of $C^{3}$, i.e., $C^{3}$ risk. Specifically, we first utilize $V$ to constrain $f_{\theta}$, effectively distinguishing $\text{F}_{c}$ from $\text{F}_{s}$ and ensuring the identifiability of $C^{3}$ based on Theorem 3.4. Obtaining $\text{Z}_{c}$ that only contains $\text{F}_{c}$, we then define sufficiency risk (when $\text{Z}=c$, the label is not $y$) and necessity risk (when $\text{Z}=\bar{c}$, the label is $y$) to construct the $C^{3}$ risk based on Definition 3.1 and Eq.5. The key challenge lies in the counterfactual data, i.e., $\overline{\text{Z}}_{c}:\text{Z}=\bar{c}$, is difficult to obtain (Kusner et al., 2017; Morgan & Winship, 2015). To address this, we propose a twin network comprising: (i) the real-world branch to obtain $\widehat{\text{Z}}_{c}$ with $V$, and (ii) the hypothetical-world branch to obtain $\overline{\text{Z}}_{c}$ under a provable gradient-based adjustment, ensuring $\overline{\text{Z}}_{c}$ stay close to the original observable distribution (Appendix D.5). These branches share network structures and parameters, maintaining a mirrored correspondence that enforces causal consistency (Pearl, 2009). Thus, we get:
Theorem 3.5 ($C^{3}$ Risk).
Given $N$ observable samples $\{(x_{i},\,y_{i})\}_{i=1}^{N}$ drawn from the multi-modal distribution $P_{XY}$, and the model $f_{\theta}$ with feature extractor $g(\cdot)$ and linear classifier $\mathcal{W}$, we define the following twin-network:
$\begin{split}&\widehat{\text{Z}}_{c,i}=g^{*}(x_{i},v_{i}),\quad\overline{\text{Z}}_{c,i}=\widehat{\text{Z}}_{c,i}+\Delta_{i},\\ &\text{where}\quad\Delta_{i}=-\nabla_{\widehat{\text{Z}}_{c,i}}\ell\Bigl{(}\sigma(\mathcal{W}^{\top}\widehat{\text{Z}}_{c,i}),\,y_{i}\Bigr{)},\\ &\text{s.t.}\quad\mu_{\text{KL}}(\overline{\text{Z}}_{c,i},\widehat{\text{Z}}_{c,i})\leq\varepsilon,\quad\forall i,\end{split}$ (7)
where $g^{*}(x_{i},v_{i})$ is the calibrated $g$ with $v_{i}$ using Kullback–Leibler divergence $\mu_{\text{KL}}(\cdot)$, i.e., with loss $\mathcal{L}_{v}=\mathbb{E}_{(x,y)\sim P_{XY}}\mathbb{E}_{v_{i}\in V}\mu_{\text{KL}}(g(x),v_{i})$, to estimate the causal representation $\widehat{\text{Z}}_{c,i}$ in the real-world branch, $\Delta_{i}$ arises from the gradient-based intervention with $v_{i}\sim V$ for the hypothetical world. If $f_{\theta}$ converges to an optimal solution under a Monte Carlo approximation of $V$, then by using:
[…]
the discrepancy between $\widehat{R}_{N}^{C^{3}}$ and the ideal causal complete quantity approaches zero with high probability. Here, $\sigma$ denotes the signum function, $\rho(\cdot)$ denotes an indicator function which is equal to 1 if the condition in $\rho(\cdot)$ is true.
Theorem 3.5 presents a method to measure the causal completeness of learned representations based on observable data. It leverages instrumental variables $V$ and a twin network to address the issues of spurious correlations in MML and the unavailability of counterfactual data in practice. The detailed proofs and analysis are provided in Appendix A.4.
4 Performance Guarantee with $C^{3}$ Risk
In this section, we conduct theoretical analyses to establish the connection between the proposed $C^{3}$ risk and MML generalization performance, proving its effectiveness.
Before discussing generalization, we first illustrate the empirical and expectation risks of the MML model $f_{\theta}$ on training and test data, i.e., $\widehat{R}^{C^{3}}(f_{\theta})$ (Eq.8) and ${R}^{C^{3}}(f_{\theta})=\mathbb{E}_{(x,y)\sim P_{XY}}\rho[\sigma(\mathcal{W}^{\top}\widehat{\text{F}}_{c,i})\neq y_{i}]+\rho[\sigma(\mathcal{W}^{\top}\overline{\text{F}}_{c,i})=y_{i}]$. Then, we provide the following performance guarantee:
Theorem 4.1 (Performance Guarantee via $C^{3}$).
Denote the hypothesis set $\mathcal{H}$ of linear classifier $\mathcal{W}\in\mathcal{H}\subseteq\{h:\mathbb{R}^{d}\rightarrow\mathcal{Y}\}$, let $H=\text{\emph{Pdim}}(\{\ell_{h}:h\in\mathcal{H}\})$. Then with probability at least $1-\delta$, and training samples that sampling from empirical distribution $\mathcal{D}_{tr}(x,y)$, we have
${R}^{C^{3}}(f_{\theta})\leq\widehat{R}^{C^{3}}(f_{\theta})+M\mathfrak{R}(\mathcal{H})+\sqrt{\frac{ln(1/\delta)}{2H}}.$ (9)
Here, $\mathfrak{R}(H)$ is the Rademacher complexity of the hypothesis class $H$, and $M$ is a finite constant.
Theorem 4.1 derives the upper bound on the generalization error of $C^{3}$ risk, relating its empirical error on the training data to its performance on unseen samples. Specifically, the generalization error is bounded by the sum of the empirical error, the Rademacher complexity of the hypothesis class of linear classifier, and a confidence term that controls the probability of exceeding the bound. This result provides a theoretical performance guarantee for optimizing MML models with $C^{3}$ risk. See Appendix A.5 for detailed proofs.
5 Learning Causal Complete Representations
Based on the above theoretical results, we propose a plug-and-play method, Causal Complete Cause Regularization ($C^{3}$), which is built upon the $C^{3}$ risk to extract causal complete representations from observable multi-modal data.
Specifically, we first introduce the constraints of the instrumental variable $V$ (Theorem 3.4) to make the MML model $f_{\theta}$ learn causal representations while satisfying the identifiability in practice. Then, we minimize the $C^{3}$ risk of the learned representations (Theorem 3.5) to ensure the causal completeness of the learned representation, i.e., causal sufficiency and necessity. In summary, the objective of $C^{3}$ is a combination of the above two-step objectives, which can be embedded in various MML models. It can be expressed as:
$\begin{split}\underset{f_{\theta}}{\min}\widehat{R}^{C^{3}}+\lambda_{v}\mathcal{L}_{v}+\lambda_{fe}\mathcal{L}_{fe}\end{split}$ (10)
where $\lambda_{v}$ and $\lambda_{fe}$ denotes the importance weights, the three terms: (i) $\widehat{R}^{C^{3}}$ (Eq.8) constrains the model to learn a causally complete representation, i.e., one that exhibits minimal sufficiency and necessity risk; (ii) $\mathcal{L}_{v}$, built upon Theorem 3.4, employs the instrumental variable $V$ to guide the model toward the causal representation, corresponding to the real-world branch; (iii) $\mathcal{L}_{fe}$ stems from the counterfactual construction condition in Eq. 7, specifically $\mathcal{L}_{fe}=\frac{1}{N}\sum_{i=1}^{N}\mu_{\text{KL}}\bigl{(}\overline{\text{Z}}_{c,i},\,\widehat{\text{Z}}_{c,i}\bigr{)}$, ensuring that $\widehat{\text{Z}}_{c,i}$ remains within the feasible region. Thus, by minimizing the above objective, the $C^{3}$ score of the learned representation will be higher while maintaining the required conditions for practical implementations. Then, it makes the learned representation causally sufficient and necessary with high confidence.
6 Experiments
In this section, we conduct extensive experiments on various benchmark datasets to verify the effectiveness of $C^{3}$. More details and experiments are provided in Appendix E-H.
Table 1: Performance comparison when 50% samples are corrupted with Gaussian noise i.e., zero mean with the variance of $N$. “(N, Avg.)” and “(N, Worst.)” denotes the average and worst-case accuracy. The best results are highlighted in bold.
Method NYU Depth V2 SUN RGB-D FOOD 101 MVSA
(0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.) (0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.) (0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.) (0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.)
CLIP (Sun et al., 2023) 69.32 68.29 51.67 48.54 56.24 54.73 35.65 32.76 85.24 84.20 52.12 49.31 62.48 61.22 31.64 28.27
ALIGN (Jia et al., 2021) 66.43 64.33 45.24 42.42 57.32 56.26 38.43 35.13 86.14 85.00 53.21 50.85 63.25 62.69 30.55 26.44
MaPLe (Khattak et al., 2023) 71.26 69.27 52.98 48.73 62.44 61.76 34.51 30.29 90.40 86.28 53.16 40.21 77.43 75.36 43.72 38.82
CoOp (Jia et al., 2022a) 67.48 66.94 49.43 45.62 58.36 56.31 39.67 35.43 88.33 85.10 55.24 51.01 74.26 73.61 42.58 37.29
VPT (Jia et al., 2022a) 62.16 61.21 41.05 37.81 54.72 53.92 33.48 29.81 83.89 82.00 51.44 49.01 65.87 64.98 32.79 29.21
Late fusion (Wang et al., 2016) 69.14 68.35 51.99 44.95 62.09 60.55 47.33 44.60 90.69 90.58 58.00 55.77 76.88 74.76 55.16 47.78
ConcatMML (Zhang et al., 2021) 70.30 69.42 53.20 47.71 61.90 61.19 45.64 42.95 89.43 88.79 56.02 54.33 75.42 75.33 53.42 50.47
AlignMML (Wang et al., 2016) 70.31 68.50 51.74 44.19 61.12 60.12 44.19 38.12 88.26 88.11 55.47 52.76 74.91 72.97 52.71 47.03
ConcatBow (Zhang et al., 2023c) 49.64 48.66 31.43 29.87 41.25 40.54 26.76 24.27 70.77 70.68 35.68 34.92 64.09 62.04 45.40 40.95
ConcatBERT (Zhang et al., 2023c) 70.56 69.83 44.52 43.29 59.76 58.92 45.85 41.76 88.20 87.81 49.86 47.79 65.59 64.74 46.12 41.81
MMTM (Joze et al., 2020) 71.04 70.18 52.28 46.18 61.72 60.94 46.03 44.28 89.75 89.43 57.91 54.98 74.24 73.55 54.63 49.72
TMC (Han et al., 2020) 71.06 69.57 53.36 49.23 60.68 60.31 45.66 41.60 89.86 89.80 61.37 61.10 74.88 71.10 60.36 53.37
LCKD (Wang et al., 2023b) 68.01 66.15 42.31 40.56 56.43 56.32 43.21 42.43 85.32 84.26 47.43 44.22 62.44 62.27 43.52 38.63
UniCODE (Xia et al., 2024) 70.12 68.74 44.78 42.79 59.21 58.55 46.32 42.21 88.39 87.21 51.28 47.95 66.97 65.94 48.34 42.95
SimMMDG (Dong et al., 2024) 71.34 70.29 45.67 44.83 60.54 60.31 47.86 45.79 89.57 88.43 52.55 50.31 67.08 66.35 49.52 44.01
MMBT (Kiela et al., 2019) 67.00 65.84 49.59 47.24 56.91 56.18 43.28 39.46 91.52 91.38 56.75 56.21 78.50 78.04 55.35 52.22
QMF (Zhang et al., 2023c) 70.09 68.81 55.60 51.07 62.09 61.30 48.58 47.50 92.92 92.72 62.21 61.76 78.07 76.30 61.28 57.61
CLIP+$C^{3}$ 76.54 75.12 56.73 52.90 62.31 58.71 41.59 37.52 92.93 91.80 59.77 57.54 69.61 68.64 39.58 35.89
MaPLe+$C^{3}$ 77.07 74.45 58.94 55.95 66.21 65.51 40.12 37.34 94.38 93.51 60.63 46.07 81.19 81.51 49.32 45.98
Late fusion+$C^{3}$ 73.26 71.62 57.21 50.98 64.84 63.25 53.35 50.43 94.09 92.24 65.27 59.02 83.77 79.79 62.14 52.50
LCKD+$C^{3}$ 77.14 75.12 50.11 47.98 60.97 60.14 47.23 46.21 90.89 90.14 54.48 51.16 66.78 65.67 49.28 42.84
SimMMDG+$C^{3}$ 75.32 74.61 49.99 47.22 65.50 64.58 52.69 51.70 92.24 91.14 57.32 53.56 73.62 71.01 51.65 51.07
MMBT+$C^{3}$ 73.74 71.82 54.35 52.57 61.47 59.99 48.42 46.07 94.25 93.90 60.41 60.11 82.76 81.64 62.12 58.93
QMF+$C^{3}$ 77.58 74.95 59.72 59.18 67.35 65.84 52.26 51.28 94.87 93.79 66.45 63.69 83.13 81.98 66.66 64.51
Table 2: Performance with missing modalities on BraTS. The brackets “()” indicate the effect changes after introducing $C^{3}$. $\bullet$” and $\circ$” indicate the availability and absence of the modality for testing. The best results are highlighted in bold.
Modalities Enhancing Tumour Tumour Core Whole Tumour
Fl T1 T1c T2 HMIS HVED RSeg mmFm LCKD LCKD+$C^{3}$ HMIS HVED RSeg mmFm LCKD LCKD+$C^{3}$ HMIS HVED RSeg mmFm LCKD LCKD+$C^{3}$
$\bullet$ $\circ$ $\circ$ $\circ$ 11.78 23.80 25.69 39.33 45.48 49.81 (+4.33) 26.06 57.90 53.57 61.21 72.01 76.65 (+4.64) 52.48 84.39 85.69 86.10 89.45 91.62 (+2.17)
$\circ$ $\bullet$ $\circ$ $\circ$ 10.16 8.60 17.29 32.53 43.22 49.13 (+6.01) 37.39 33.90 47.90 56.55 66.58 72.18 (+5.60) 57.62 49.51 70.11 67.52 76.48 82.39 (+5.91)
$\circ$ $\circ$ $\bullet$ $\circ$ 62.02 57.64 67.07 72.60 75.65 80.50 (+4.85) 65.29 59.59 76.83 75.41 83.02 88.06 (+5.04) 61.53 53.62 73.31 72.22 77.23 81.93 (+4.70)
$\circ$ $\circ$ $\circ$ $\bullet$ 25.63 22.82 28.97 43.05 47.19 54.13 (+6.94) 57.20 54.67 57.49 64.20 70.17 77.32 (+7.15) 80.96 79.83 82.24 81.15 84.37 90.78 (+6.41)
$\bullet$ $\bullet$ $\circ$ $\circ$ 10.71 27.96 32.13 42.96 48.30 54.16 (+4.86) 41.12 61.14 60.68 65.91 74.58 79.83 (+5.25) 64.62 85.71 88.24 87.06 89.97 93.63 (+3.66)
$\bullet$ $\circ$ $\bullet$ $\circ$ 66.10 68.36 70.30 75.07 78.75 82.98 (+4.23) 71.49 75.07 80.62 77.88 85.67 89.74 (+4.07) 68.99 85.93 88.51 87.30 90.47 93.91 (+3.44)
$\bullet$ $\circ$ $\circ$ $\bullet$ 30.22 32.31 33.84 47.52 49.01 56.12 (+7.11) 57.68 62.70 61.16 69.75 75.41 82.57 (+7.16) 82.95 87.58 88.28 87.59 90.39 95.48 (+5.09)
$\circ$ $\bullet$ $\bullet$ $\circ$ 66.22 61.11 69.06 74.04 76.09 81.76 (+5.67) 72.46 67.55 78.72 78.59 82.49 88.32 (+5.83) 68.47 64.22 77.18 74.42 80.10 87.03 (+6.96)
$\circ$ $\bullet$ $\circ$ $\bullet$ 32.39 24.29 32.01 44.99 50.09 56.03 (+5.94) 60.92 56.26 62.19 69.42 72.75 78.78 (+6.03) 82.41 81.56 84.78 82.20 86.05 92.33 (+6.28)
$\circ$ $\circ$ $\bullet$ $\bullet$ 67.83 67.83 69.71 74.51 76.01 83.97 (+7.96) 76.64 73.92 80.20 78.61 84.85 93.57 (+8.72) 82.48 81.32 85.19 82.99 86.49 94.40 (+7.91)
$\bullet$ $\bullet$ $\bullet$ $\circ$ 68.54 68.60 70.78 75.47 77.78 82.94 (+5.06) 76.01 77.05 81.06 79.80 85.24 90.46 (+5.22) 72.31 86.72 88.73 87.33 90.50 95.51 (+5.01)
$\bullet$ $\bullet$ $\circ$ $\bullet$ 31.07 32.34 36.41 47.70 49.96 56.25 (+6.29) 60.32 63.14 64.38 71.52 76.68 82.69 (+6.01) 83.43 88.07 88.81 87.75 90.46 96.23 (+5.77)
$\bullet$ $\circ$ $\bullet$ $\bullet$ 68.72 68.93 70.88 75.67 77.48 83.90 (+6.42) 77.53 76.75 80.72 79.55 85.56 92.39 (+6.83) 83.85 88.09 89.27 88.14 90.90 96.78 (+5.88)
$\circ$ $\bullet$ $\bullet$ $\bullet$ 69.92 67.75 70.10 74.75 77.60 82.54 (+4.94) 78.96 75.28 80.33 80.39 84.02 89.43 (+5.41) 83.94 82.32 86.01 82.71 86.73 91.73 (+5.00)
$\bullet$ $\bullet$ $\bullet$ $\bullet$ 70.24 69.03 71.13 77.61 79.33 86.36 (+7.03) 79.48 77.71 80.86 85.78 85.31 91.43 (+6.12) 84.74 88.46 89.45 89.64 90.84 95.41 (+4.57)
Figure 3: Evaluation for the property of learned representations (SNC, SC, NC, and SP). See Appendix H for details.
Figure 4: Ablation study of $C^{3}$ (performance when removing different regular terms). See Appendix H for details.
6.1 Experimental Settings
Datasets We select six datasets: (i) scenes recognition on NYU Depth V2 (Silberman et al., 2012) and SUN RGBD (Song et al., 2015) with RGB and depth images; (ii) image-text classification on UPMC FOOD101 (Wang et al., 2015) and MVSA (Niu et al., 2016) with image and text; (iii) segmentation considering missing modalities on BraTS (Menze et al., 2014; Bakas et al., 2018) with Flair, T1, T1c, and T2; and (iv) synthetic MMLSynData (see Appendix D.1).
Implementation Details We use a three-layer MLP with activation functions (Clevert et al., 2015) as the representation learner. The hidden vector dimensions of each layer are specified as 64, 32, and 128, while the learned representation is 64. For optimization, we employ the Adam optimizer (Kingma & Ba, 2014) with Momentum and weight decay set at $0.8$ and $10^{-4}$. The initial learning rate is established at 0.1, with the flexibility for linear scaling as required. Additionally, we use grid search to set the hyperparameters $\lambda_{v}=0.75$ and $\lambda_{fe}=0.4$. All experimental procedures are executed in five runs via NVIDIA RTX A6000 GPUs.
6.2 Results
Performance and robustness analysis
To evaluate the effectiveness of $C^{3}$, we record the average and worst-case accuracy of various MML baselines and introduce $C^{3}$ under Gaussian noise (for image modality) and blank noise (for text modality) following (Han et al., 2020; Zhang et al., 2023d; Ma et al., 2021). The results are shown in Table 1. We can observe that $C^{3}$ achieves stable improvements in both the average and worst-case accuracy. This proves the superior effectiveness and robustness of $C^{3}$.
When faces the problem of missing modalities
Considering the missing modality issues, we evaluate the performance of $C^{3}$ and several strong baselines (Wang et al., 2023b; Bakas et al., 2018; Zhang et al., 2022) on all 15 possible combinations of missing modalities on BraTS. From Table 2, we can observe that (i) $C^{3}$ brings significant performance improvements; (ii) $C^{3}$ can reduce the learning gap for the representations on different modal semantics, i.e., reducing the accuracy gap of learning on the difficult-to-identify Fl and T1 modalities and the easy T1c. This demonstrates the superiority of $C^{3}$ and the advantage of causally complete representation in missing modality issues.
Learning causal complete representations
To evaluate $C^{3}$’s ability to extract causal complete causes, we conduct experiments: (i) construct four types of MML data, i.e., sufficient and necessary (SNC), sufficient but unnecessary (SC), necessary but insufficient causes (NC), and spurious correlations (SP) following (Yang et al., 2024) (see Appendix H for more details); then (ii) evaluate their correlation with the learned representation based on the distance metric (Jones et al., 1995), i.e., the higher the score, the stronger the correlation. The results are shown in Figure 4 (see Appendix D and H for details). The representations learned by $C^{3}$ have a higher correlation with SNC and lower with SP. This proves that $C^{3}$ can learn causal complete representations more effectively while other methods are hard to achieve.
Ablation Study
To evaluate the effect of each item in $C^{3}$ (Eq.10), we evaluate the performance change on LCKD+$C^{3}$. We consider the standard and corner case (missing modality), and conduct experiments on FOOD 101 and BraTS, as shown in Figure 4. We can observe that each item plays an important role. See Appendix H for more experiments, e.g., model efficiency and parameter sensitivity.
7 Related Work
Multi-modal learning aims to learn good representations through multiple modalities for accurate prediction. Recently, multiple methods (Xu et al., 2023; Jia et al., 2022b; Wang et al., 2023c; Fan et al., 2024) have been proposed to solve MML tasks through modality consistency, e.g., tokenize diverse modalities into sequences and utilize Transformers for joint learning (Bao et al., 2022; Wang et al., 2022), whereas CLIP (Radford et al., 2021; Fan et al., 2024), ALIGN (Jia et al., 2021), etc. employ distinct encoders for each modality and utilize contrastive loss to synchronize features. These methods align features from different modalities into the same space to capture the primary events. Another type of work (Jiang et al., 2023; Dong et al., 2024; Wang et al., 2023a; Wu et al., 2020) proposed to additionally focus on modality specificity. They divide the MML features into modality-shared and modality-specific components and learn separately. However, all these methods may result in learning insufficient or unnecessary information. Meanwhile, they rely on strong assumptions like semantic alignment (Akbari et al., 2021; Wang et al., 2023c; Lu et al., 2022; Zhang et al., 2023b) and may limited to ideal data (Ge et al., 2023; Zhang et al., 2023a), affecting model performance. For causal sufficiency and necessity, (Pearl, 2009) first proposed the related concepts, and (Yang et al., 2024) applied it to domain generalization with assumptions in an ideal environment. The differences between our methodology and theirs include problem settings, theorems, implementations, learning objectives, etc. (See Appendix D.6 for detailed comparison). We explore the MML-specific concepts of causal completeness without strong assumptions, ensuring the effectiveness of the learned representations.
8 Conclusion
We explore the MML-specific concepts of causal sufficient and necessary without exogeneity and monotonicity assumptions. To measure whether the learned MML representation is causally complete, we present the definition, identifiability, and measurement of $C^{3}$ with theoretical support. Based on these results, we propose a plug-and-play method $C^{3}$ to promote MML learning causal complete representations. Extensive experiments demonstrate its effectiveness.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and Multi-Modal Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.